My daughter has been struggling tonight with this algebra problem and I couldn't help! Please help! "Six times as many grandstand tickets g as box-seat tickets b were sold for the baseball game. Grandstand tickets cost $10 and box-seat tickets cost $25. The total amount of ticket sales for grandstand and box-seat tickets was $67,320. We need to write a system of equations that describes the above situation and solve for the number of box seats. We believe one equation is 10g + 25b = $67,320, but don't know where to go from there. Help.

Answer by scott8148(5891)

good so far ... your missing equation is 6b=g (see first sentence of problem)

substituting into the equation you found gives 10(6b)+25b=67320 ... 60b+25b=67320 ...85b=67320 ... b=792 ... g=4752